# David Roodman Open Philanthropy San Francisco, CA

### v2

@@ 4.3 Limitations on accessing Stata objects
You are not supposed to call the plugin directly. But you could. For example:
  • plugin call _julia, eval "1+1"
The designated entry point to the plugin, the C++ function stata_call(), would then receive and process the two arguments after the comma, which tell it to have Julia evaluate “1+1”. Since calls to the plugin are in general arcane and sometimes require preparatory steps, the jl program serves as an essential and more user-friendly wrapper. That extra layer complicates access to local macros. If Julia, having been invoked by jl and the plugin, uses the SPI function SF_macro_save() to write a Stata local named foo, then foo will appear in the macro name space of the program that called the plugin, which is jl. It will not appear in the context a step up, which could be your code. As a workaround, every time the Julia function st_local() is called, it adds the name of the macro it is writing to another local, __jllocals. Before exiting, jl copies every macro in that list to its caller’s context using the undocumented Stata command c_local.
However, the same complication prevents one from reading Stata locals from Julia. The workaround just described only goes in one direction, because I know of no way for an ado program to read its caller’s local macros. In partial compensation, single-line jl: commands can quote locals, because Stata processes the lines before passing them to jl. For example:
@@
This trick will not work in an interactive Julia session.
@@ 4.4 Speed considerations in copying data: type conversions, missingness, and subsample restrictions
The specification of the Stata plugin interface introduces several wrinkles into the copying of data between Stata and Julia.
One is that the functions for reading and writing numeric Stata variables, SF_vdata() and SF_vstore(), only work with double-precision values, regardless of whether a Stata-side variable is stored as byte, int, long, float, or double. By default, the jl commands that copy from the active Stata data set to Julia DataFrames—jl save and jl PutVarsToDF—convert the double-precision values back to the original data types. [^13] Because the ranges of Stata’s byte, int, and long types are narrowed in the way described in the next paragraph, when transferring data from Julia to Stata, the corresponding Julia types are mapped to larger types in Stata. For example, columns of type Int8, which have an allowed range of $[-128,+127]$, are copied to Stata’s two-byte int type, since Stata bytes only hold $[-128,100]$. So a Stata byte variable will be converted to double and back on the way to appearing in a Julia DataFrame. The jl commands can do nothing about the first conversion. But they do accept a doubleonly option to skip the second conversion and instead leave all copied values in double precision. This typically saves time, both by reducing conversions and by allowing Julia to allocate a contiguous block of memory to receive all the data, rather than a separate block for each of variously typed columns.
Separately, Stata and Julia represent missing values in different ways. Stata reserves certain high values of each data type for 27 flavors of missing: . and .a, …, .z. That is why the maximum value of a Stata byte is 100, not than 127. Julia only supports one flavor of missing, and stores missingness information alongside rather than in data values. It also allows observations of string variables to be missing. The julia package translates missings when copying between Stata data sets and Julia DataFrames (but not Julia matrices). For data known to contain no missing values, adding the nomissing option will save time by skipping this step.
Finally, the SPI lets plugin calls include if or in clauses to restrict the sample on the Stata side. The julia plugin contains optimized code that runs when such clauses are not specified.
@@ 4.5 Simulated read-eval-print loop
The Julia language system has many components—for interpreting and compiling code, managing memory, and so on. The core parts are written in C++ and distributed as compiled libraries. Other components are themselves written in Julia, including the front end, which users often think of as Julia per se. The front end is usually called the REPL, for “Read-Eval-Print Loop.” Indeed, the REPL’s core job is to read in commands, evaluate them, print the results, and repeat. But the REPL provides other conveniences: typing ”?” starts help mode; typing ”]” instead switches to the Julia package manager; some use is made of color; arrow keys allow one to explore one’s command history.
While there may be a way to run the full Julia REPL inside of Stata, the jl command does not do that. It gins up a limited REPL. That is because The plugin directly accesses lower-level components of Julia, through the C-language entry points that loosely comprise Julia’s interface for “embedding” in C. A workhorse for jl is the Julia system’s jl_eval_string() function. It accepts a C string holding a Julia expression, evaluates it, and returns the result in a C structure that represents a generic Julia object. All Julia lines submitted to jl pass through this function. As its name suggests, the job of the function is to evaluate strings, not provide a user experience.
@@
    exit()
This snippet prints the sum of the first 10 whole numbers—at least when run in the Julia REPL. Actually, for reasons having nothing to do with Stata, the snippet will crash if run non-interactively, as part of a stored program. Under Julia’s “soft scoping” rule, since S is not declared or initialized within the for loop before the first attempt to increment it, the attempt will trigger an undefined variable error. Because that rule is counterintuitive, Julia’s REPL engineers an exception for interactive use. [^14] See the documentation at docs.julialang.org/en/v1/manual/variables-and-scoping/#local-scope. jl: imitates the exception by calling the Julia softscope() function on the user’s input after parsing it. softscope() is itself part of the code base of the Julia REPL.
@@
And the processing required for the simulation, along with the special handling of Stata locals, adds time to the execution of each Julia line from Stata. The extra 0.02 seconds or so is not noticeable in an interactive session. But it can add up in a program that calls jl: many times. The _jl: variant of the prefix command bypasses all of this processing.
@@ 5.1 Architecture
The julia package consists of three intertwined pieces, all of which have been mentioned. The core is a shared library written in C++ according to the specification for Stata plugins. It comes precompiled for Windows, Linux, and macOS with Intel or Apple silicon. The “julia.pkg” manifest file instructs Stata to select the copy of the library that will run on the target computer, and to name it “jl.plugin.” The plugin performs two main jobs: efficiently copying data among matrices and data sets in Julia and Stata, and sending commands to Julia and returning the results.
The file “jl.ado” serves as a Stata front end for the plugin, providing a more intuitive interface and performing tasks better done in the ado language than in C++. At the suggestion of Jeffrey Pitblado at StataCorp, this program is called jl rather than julia because Stata might introduce an official julia command someday. When jl is first called in a Stata session, it briefly runs the Julia REPL through a shell command in order to query the location on disk of Julia’s shared libraries. It passes the location to the plugin, which loads the libraries. This lets the plugin access Julia.
jl also calls Stata’s findfile to track down the third key piece of the package. “stataplugininterface.jl” defines a module of Stata interface functions in Julia, such as st_local() and st_data(). These work by calling entry points in the Stata plugin interface, which are exported to Julia by the plugin. (The Julia language includes a way to call C functions.) At start-up, jl therefore stores the location of “jl.plugin” in the stataplugininterface module.
@@ 5.2 Stata commands in the julia package
jl also accepts subcommands. As discussed in section 4.2, one is especially useful in macOS, because it allows users to set the upper limit on the number of threads:
@@
Several subcommands help manage Julia packages. Just as Stata downloads packages into folders such as “/home/droodman/ado/plus,” Julia downloads and stores packages in directories called environments. Julia also makes it easy to change the current environment. This is useful for the developer of a Julia-calling Stata program because it allows the Stata program to install needed Julia packages without interfering with the user’s package configuration. The package management subcommands are:
@@
GetEnv displays the directory of the current package environment and returns it in r() macros. On my computer, which is running Julia 1.10, the command displays:
  • . jl GetEnv
[…]
jl SetEnv moves to a chosen, named environment, which is merely a subdirectory of the default environment directory:
@@
If the environment does not exist, it is created and populated with the DataFrames.jl and CategoricalArrays.jl packages. One can return to the default environment with SetEnv . or just SetEnv. [^15] In distributed computing, Julia worker processes do not automatically inherit the master’s package environment. One can pass on the master’s environment through command line switches when launching the workers, such as in addprocs(4, exeflags="--project=\$(Base.active_project())").
Last in the package management family is the AddPkg option, which installs a package in the current environment or, if minver() is specified, updates it to the latest version if the current version falls below the minimum. Versions X.Y.Z are in the three-part semantic versioning format.
The remaining jl subcommands copy data between Stata and Julia. The first six listed below are meant more as programmer’s commands while the last two are higher-level. jl GetVarsFromDF, for example, requires that the Stata data set already be at least as tall as the DataFrame from which it will receive variables. In contrast, jl use works like Stata’s use command: it replaces the current data set, and will fail unless the replace option is included, or the current data set has not changed since it was last saved.
  • jl PutVarsToDF [varlist] [if] [in], [destination(string) cols(string) nolabel nomissing doubleonly] jl GetVarsFromDF varlist [if] [in], [cols(string) source(string) replace nomissing] jl PutVarsToMat [varlist] [if] [in], destination(string) jl GetVarsFromMat varlist [if] [in], source(string) jl PutMatToMat matname, [destination(string)] jl GetMatFromMat matname, [source(string)] jl save [dataframename], [nolabel nomissing doubleonly] jl use dataframename, [clear] jl use varlist using dataframename, [clear]
The varlists and matnames before the commas refer to Stata variables or matrices. If an optional varlist is omitted, it defaults to *, meaning all variables in the current data set in their current order. A destination() or source() option after the comma refers to a Julia matrix or DataFrame. When an optional DataFrame name is not provided, it defaults to df. The cols() option specifies the DataFrame columns to be copied to or from. It defaults to the Stata varlist before the comma. Destination Stata matrices and Julia matrices and DataFrames are entirely replaced. The low-level commands will create the destination Stata variables or, if replace is specified, overwrite them subject to any [if] or [in] restriction.
@@ 5.3 Julia functions in the julia package
The full list of functions:
@@
SF_var_is_binary(i, j) Whether observation i of variable j is a binary strL
@@
SV_missval() Stata floating-point value for missing
@@
SF_macro_use(mac, maxlen) First maxlen characters of macro mac
@@
st_matrix(matname, jlmat) Put Julia matrix in pre-existing Stata matrix
@@
      v = st_view("price mpg", sample)  # view onto subsample of two vars
      v ./= 2                           # halve these Stata data points
@@ 6 Conclusions
Julia is a strong option for universal, back-end development of statistical applications. Working in Julia comes with some drawbacks, including the challenge of learning a complex new language, the delays upon installation and first use of programs, and the immaturity of some packages and their documentation. But Julia and nearly all Julia packages are free and open and run on all popular operating systems. The language blends features for maximizing performance where it matters with a Python-like expressiveness that makes code easier to write and maintain. The package ecosystem covers domains essential to efficient statistical work, such as multiprocessing, linear algebra, and numerical optimization.
Instead of implementing new methods separately in Stata or R or Python, or in a low-level language such as C, developers can write the back ends once, in Julia. Here, WildBootTests.jl shows the way: it is an optional back end for boottest in Stata and a required back end for instrumental variables inference with fwildclusterboot in R.
For Stata users and developers, the julia package demonstrates the practicality and value of bridging from Stata to Julia. Through a plugin, data can be piped at high speed. Most features needed to make Stata objects accessible in Julia can and have been implemented. The linkage gives Stata users access to Julia packages for estimation and plotting that offer greater speed, new features, or more flexible syntax (as for plotting). reghdfejl is a perhaps the “killer app” for the Julia-Stata link. It fits linear models much faster than the pioneering reghdfe, at least for problems that are time consuming for reghdfe. And reghdfejl preliminarily handles nonlinear HDFE models, which brings new functionality to Stata users.
StataCorp could strengthen Stata’s bond to Julia, as it has done for Python and Java. Stata’s developers could, for example, extend the plugin interface to allow creation of Stata matrices and variables. They could add a Julia mode to the do file processor, so that Julia code could be embedded without awkward semicolons and line continuation symbols. They might be able to grant more access to Stata variable storage en bloc, so that Julia code could read and write Stata data without copying it first. They could make Mata interoperate with Julia.

### v3

@@ 4.3 Limitations in accessing Stata objects
You are not supposed to call the plugin directly. But after typing that line, you could. For example:
  • plugin call _julia, eval "x=1"
The designated entry point to the plugin, the C++ function stata_call(), would then receive and process the two arguments after the comma, which tell it to have Julia evaluate x=1. Since calls to the plugin are in general arcane and sometimes require preparatory steps, the jl program serves as an essential and more user-friendly wrapper for it. That extra layer complicates access to local macros. If Julia, having been invoked by jl and the plugin, uses the SPI function SF_macro_save() to write a Stata local named foo, then foo will appear in the macro name space of the program that called the plugin, which is jl. It will not appear in the context a step up, which could be your code. As a workaround, every time the Julia function st_local() is called, it adds the name of the macro it is writing to another local in the name space of the jl program, called __jllocals. Before exiting, jl copies every macro in that list to its caller’s context using the undocumented Stata command c_local.
However, the same complication prevents one from reading Stata locals from Julia. The workaround just described only goes in one direction, because I know of no way for an ado program to read its caller’s local macros.
As partial compensation, single-line jl: commands can quote locals, because Stata processes the lines before passing them to jl. For example:
@@
This trick will not work in an interactive Julia session. [^14] The single, compound line jl: st_local("a","123"); print("‘a’") will also fail. The second command’s reference to a will be expanded before control is passed to Julia. The commands will work as intended if split into two lines.
@@ 4.4 Speed considerations in copying data: type conversions, missingness, and subsample restrictions
StataCorp’s definition of the Stata plugin interface introduces several wrinkles into the copying of data between Stata and Julia.
One is that the functions for reading and writing numeric Stata variables, SF_vdata() and SF_vstore(), only work with double-precision values, regardless of whether a Stata-side variable is stored as byte, int, long, float, or double. By default, the jl commands that copy from the active Stata data set to Julia DataFrames—jl save and jl PutVarsToDF—convert the double-precision values back to the original data types. [^15] Because the ranges of Stata’s byte, int, and long types are narrowed in the way described in the next paragraph, when transferring data from Julia to Stata, the corresponding Julia types are mapped to larger types in Stata. For example, columns of type Int8, which have an allowed range of $[-128,+127]$, are copied to Stata’s two-byte int type, since Stata bytes only hold $[-128,100]$. So a Stata byte variable will be converted to double and back on the way to appearing in a Julia DataFrame. The jl commands can do nothing about the first conversion since that is performed by Stata. But they do accept a doubleonly option to skip the second conversion and instead leave all copied values in double precision. This typically saves time, both by reducing conversions and by allowing Julia to allocate a contiguous block of memory to receive purely double-precision values, rather than a separate block for each of variously typed columns.
Separately, Stata and Julia represent missing values in different ways. Stata reserves certain high values of each data type for 27 flavors of missing: . and .a, …, .z. That is why the maximum value of a Stata byte is 100, not than 127. Julia only supports one flavor of missing, and stores missingness information alongside rather than in data values. It thus distinguishes between empty and missing strings. The jl subcommands translate missings when copying between Stata data sets and Julia DataFrames (but not Julia matrices). For data known to contain no missing values, adding the nomissing option will save time by skipping this step.
Finally, the SPI lets plugin calls include if or in clauses to restrict the sample on the Stata side. When such clauses are not specified, the julia plugin runs specially optimized code.
@@ 4.5 Simulated read-eval-print loop
The Julia language system has many components—for compiling code, managing memory, and so on. The core parts are written in C++ and distributed as compiled libraries. Other components are themselves written in Julia, including the front end, which users often think of as Julia per se. The front end is called the REPL, for “Read-Eval-Print Loop.” Indeed, the REPL’s core job is to read in commands, evaluate them, print the results, and repeat. But the REPL provides other conveniences. Typing ”?” starts help mode. Typing ”]” instead switches to the Julia package manager. Greek letters and other symbols can be entered through a TeX-like syntax. Arrow keys allow one to explore one’s command history.
While there may be a way to run the full Julia REPL inside of Stata, the jl command does not do that. It gins up a limited REPL. That is because the plugin directly accesses lower-level components of Julia, through the C-language entry points that loosely comprise Julia’s interface for “embedding” in C. A workhorse for jl is the Julia system’s jl_eval_string() function. It accepts a C string holding a Julia expression, evaluates it, and returns the result in a C data structure that represents a generic Julia object. All Julia lines submitted to jl pass through this function. As its name suggests, the function’s job is to evaluate strings, not provide a user experience.
@@
This snippet prints the sum of the first 10 whole numbers—at least when run in the Julia REPL. Actually, for reasons having nothing to do with Stata, the snippet will crash if run non-interactively, as part of a stored Julia program file. Under Julia’s “soft scoping” rule, since S is declared as a global variable outside the for loop or any other structure such as a function definition, it is invisible within the for loop. The first attempt to increment it from within the loop will trigger an undefined-variable error. Because that rule is counterintuitive, Julia’s REPL engineers an exception for interactive use. [^16] See the documentation at docs.julialang.org/en/v1/manual/variables-and-scoping/#local-scope. jl imitates the exception by calling the Julia softscope() function on the user’s input after parsing it. softscope() is itself part of the code base of the Julia REPL. [^17] Thanks to Mark Kittisopikul for the idea of calling softscope().
@@
And the processing required for the simulation, along with the special handling of Stata locals, adds time to the execution of each Julia line from Stata. The extra 0.02 seconds or so is not noticeable in an interactive session. But it can add up in a program that calls jl: many times. The _jl: variant of the prefix command bypasses all of this REPL simulation.
@@ 5.1 Architecture
The julia package consists of three intertwined pieces. The core is a shared library written in C++ according to the specification for Stata plugins. It comes precompiled for Windows, Linux, and macOS with Intel or Apple silicon. The “julia.pkg” manifest file instructs Stata to select the copy of the library that will run on the target computer, and to name it “jl.plugin.” The plugin performs two main jobs: efficiently copying data among matrices and data sets in Julia and Stata, and sending user-written commands to Julia and returning the results.
The file “jl.ado” serves as a Stata front end for the plugin, providing a more intuitive interface and performing tasks better done in the ado language than in C++. At the suggestion of Jeffrey Pitblado at StataCorp, this program is called jl rather than julia because Stata might introduce an official julia command. When jl is first called in a Stata session, it briefly runs the true Julia REPL through a shell command in order to query the location on disk of Julia’s shared libraries. It passes the location to the plugin, which loads the libraries. [^18] Thanks to David Anthoff and Elliot Saba for guidance on this solution.
jl also calls Stata’s findfile to track down the third key piece of the package. “stataplugininterface.jl” defines a module of Stata interface functions in Julia, such as st_numscalar() and st_matrix(), which are used in examples above. [^19] It would be more “Julianic” to name the file “StataPluginInterface.jl”. But rather than being posted to Julia’s general package registry, the file is distributed as part of a Stata package, and the Stata package installer cannot handle uppercase characters. These work by calling C-language entry points in the Stata plugin interface, which are exported to Julia by the plugin. (The Julia language includes a way to call C functions.) At start-up, jl therefore stores the location of “jl.plugin” in a variable in the loaded stataplugininterface module.
@@ 5.2 Stata commands in the julia package
jl also accepts subcommands. As discussed in section 4.2, one is especially useful in macOS, because it allows users to cap the number of threads:
@@
Several subcommands help manage Julia packages, which are discrete libraries of code structured for easy sharing and are usually hosted on GitHub. Just as Stata downloads packages into folders such as “/home/droodman/ado/plus,” Julia downloads and stores packages in directories called environments. Julia also makes it easy to change the current environment. This is useful for the developer of a Julia-calling Stata program because it allows the Stata program to install needed Julia packages without interfering with the user’s package configuration. The package management subcommands are:
@@
GetEnv displays the directory and contents of the current package environment and returns them in r() macros. jl SetEnv moves to a chosen, named environment, which is merely a subdirectory of the default environment directory. If the environment does not exist, it is created and populated with the DataFrames.jl and CategoricalArrays.jl packages. One can return to the default environment with “SetEnv .” or just “SetEnv”. [^20] In distributed computing, Julia worker processes do not automatically inherit the master’s package environment. One can pass on the master’s environment through command line switches when launching the workers, such as in addprocs(4, exeflags="--project=\$(Base.active_project())").
On my computer, creating a new environment called myenv produces this output:
@@
[…]
      [324d7699] CategoricalArrays v0.10.8
      [a93c6f00] DataFrames v1.6.1
The output ends with a list of currently installed packages—here, the two that are automatically installed.
Last in the package management family is the AddPkg subcommand. Within the current environment, the subcommand installs a package; or, if minver() is specified, updates the package to the latest version if the current version falls below the specified minimum. Versions X.Y.Z are in the three-part semantic versioning format.
The remaining jl subcommands copy data between Stata and Julia. The first six listed below are meant more as programmer’s commands while the last two are higher-level. [^21] Johannes Boehm suggested adding high-level data-copying commands. jl GetVarsFromDF, for example, requires that the Stata data set already be at least as tall as the DataFrame from which it will receive variables. In contrast, jl use works like Stata’s use command: it replaces the current data set, and will fail unless the replace option is included, or the current data set has not changed since it was last saved.
  • jl PutVarsToDF [varlist] [if] [in], [destination(string) cols(string) nolabel nomissing doubleonly] jl GetVarsFromDF varlist [if] [in], [cols(string) source(string) replace nomissing] jl PutVarsToMat [varlist] [if] [in], destination(string) jl GetVarsFromMat varlist [if] [in], source(string) [replace] jl PutMatToMat matname, [destination(string)] jl GetMatFromMat matname, [source(string)] jl save [dataframename], [nolabel nomissing doubleonly] jl use dataframename, [clear] jl use varlist using dataframename, [clear]
The varlists and matnames before the commas refer to Stata variables or matrices. If an optional varlist is omitted, it defaults to *, meaning all variables in the current data set in their current order. A destination() or source() option after the comma should contain the name of a Julia matrix or DataFrame. When an optional DataFrame name is not provided, it defaults to df. The cols() option specifies the DataFrame columns to be copied to or from. It defaults to the Stata varlist before the comma. Destination Stata matrices and Julia matrices and DataFrames are entirely replaced. The low-level commands will create the destination Stata variables or, if replace is specified, overwrite them subject to any [if] or [in] restriction.
@@ 5.3 Julia functions in the julia package
The full list of Stata-interfacing Julia functions:
@@
SF_var_is_binary(j, i) Whether observation j of variable i is a binary strL
@@
SF_missval() Stata floating-point value for missing
@@
SF_macro_use(mac) Contents of macro mac
@@
st_matrix(matname, jlmat) Put Julia matrix in existing Stata matrix
@@
      v = st_view("price mpg", sample)  # view onto subsample of two Stata variables
      v ./= 2                           # halve these values in the Stata data set
@@ 6 Conclusions
Julia is a strong option for universal back-end development of statistical applications. Working in Julia comes with some drawbacks, including the challenge of learning a complex language, the delays upon installation and first use of programs, and the immaturity of some packages and their documentation. But Julia and nearly all Julia packages are free and open and run on all popular operating systems. The language blends features for maximizing performance where it matters with a Python-like expressiveness that makes code easier to write and maintain. The package ecosystem covers domains essential to efficient statistical work, such as multiprocessing, linear algebra, and numerical optimization.
Instead of implementing new methods separately in Stata or R or Python, or in a low-level language such as C, developers can write the back ends once, in Julia. Here, WildBootTests.jl shows the way: it is an optional back end for boottest in Stata and fwildclusterboot in R, and a required back end for instrumental variables inference with the latter.
For Stata users and developers, the julia package demonstrates the practicality and value of bridging from Stata to Julia. Through a plugin, data can be piped at high speed. Most features needed to make Stata objects accessible in Julia can and have been implemented. The linkage gives Stata users access to Julia packages for estimation and plotting that offer greater speed or new features. reghdfejl might be the “killer app” for the Julia-Stata link. It fits linear models much faster than the pioneering reghdfe, at least for problems that are time consuming for the older program. It also handily outpaces the newly augmented areg. And, preliminarily, reghdfejl handles nonlinear HDFE models, thus bringing new functionality to Stata users.
StataCorp could strengthen Stata’s bond to Julia, as it has done for Python and Java. Stata’s developers could, for example, extend the plugin interface to allow creation of Stata matrices and variables. They could add a Julia mode to the do file processor, so that Julia code could be embedded without awkward semicolons and line continuation symbols. They might be able to grant more access to Stata variable storage en bloc, so that Julia code could read and write Stata data without copying it one observation at a time, or even without copying it at all. They could give Mata the ability to call Julia functions.
@@
7 Acknowledgments
I thank Sergio Correia, Páll Haraldsson, and an anonymous reviewer for comments.